# PointNetPGAP-SLC: A 3D LiDAR-based Place Recognition Approach with Segment-level Consistency Training for Mobile Robots in Horticulture

### v2

@@ Abstract
00footnotetext: 1T.Barros, L.Garrote, P.Conde, C.Premebida and U.J.Nunes are with the University of Coimbra, Institute of Systems and Robotics, Department of Electrical and Computer Engineering, Portugal. E-mails: {tiagobarros, garrote, pedro.conde,  cpremebida, urbano}@isr.uc.pt00footnotetext: 2 M.J.Coombes and C.Liu are with the Dept. of Aeronautical and Automotive Engineering, LUCAS Lab, Loughborough University, UK. E-mails:{M.J.Coombes, C.Liu5}@lboro.ac.uk
@@ I INTRODUCTION
In recent years, significant advancements have been made in 3D LiDAR-based place recognition, driven primarily by the autonomous driving community with the aim of achieving long-term localization [1, 2]. While initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved. Unlike urban environments, where LiDAR scans capture well-defined geometries (e.g., corners, planes), horticultural environments produce scans with sparse and poorly defined features. The semi-permeable nature of these environments allows laser beams to traverse multiple row layers, resulting in scans that cover several rows. This phenomenon leads to sparse scans that overlap with scans from adjacent rows, thereby causing both intra- and inter-row descriptor ambiguity.
@@
We benchmark PointNetPGAP at both the sequence and segment levels against state-of-the-art (SOTA) models, specifically OverlapTransformer [5], PointNetVLAD [6], and LOGG3D-Net [7], using the HORTO-3DLM dataset within a cross-validation protocol. Additionally, we evaluate the cross-domain generalization capabilities using the KITTI Odometry dataset [8].
The results highlight that the aggregation approach of PointNetPGAP is better suited to horticultural environments compared to existing SOTA models. Although its impact is reduced in urban settings, it nonetheless exhibits superior rotation invariance in urban environments compared to SOTA models. Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved , particularly by improving retrieval accuracy in segments with higher ambiguity.
@@ II RELATED WORK
In the context of 3D LiDAR-based methods, although hand-crafted descriptors are still being used [9], Deep Learning (DL) techniques have emerged as the most prevalent choice for place modeling [1]. DL-based place recognition models can be broadly split into two main components: a local feature extraction module that maps the input scan to a local feature space, and a global feature aggregation module that aggregates the local features into a global descriptor.
As for the feature extraction module, approaches can be categorized based on how the point clouds are processed: those that extract features directly from point clouds, such as PointNetVLAD [6], and those that resort to proxy representations, such as voxels [10, 11], polar coordinates [12], or depth range images [13, 5]. Recent works such as MinkLoc3D [11], LCDNet [14], and LOGG3D-Net [7] tend to primarily use voxel-based approaches, employing sparse convolutions such as PV-RCNN [15] or 4D Spatio-Temporal ConvNets [16], which have shown higher feature extraction capabilities in urban-like environments. In this work, we resort to PointNet [17] due to its efficiency and lower computational demands (as demonstrated in Section IV-F).
Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [6, 18, 14], which computes the residuals w.r.t. feature clusters, generalized-mean pooling (GeM) [11] or global max pooling (MAC) [19] which operate directly on the feature space. Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach. Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor. We compare our hypothesis with the SOTA models LOGG3D-Net [7], PointNetVLAD [6], and OverlapTransformer [5], and show through results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
@@ III Proposed Approach
This section introduces the PointNetPGAP and the proposed training regime, where a traditional contrastive learning approach is combined with the proposed SLC model. Figure 2 outlines both PointNetPGAP in a retrieval framework (Top) and PointNetPGAP in the proposed training scheme (Bottom), with the combination of both the contrastive and the Negative Log-Likelihood (NLL) losses.
@@ III-B PointNetPGAP
PointNetPGAP, here defined as a mapping function $\Theta$ (outlined in Fig. 2 by T1, T2), can be decomposed into two functions: $\Theta\equiv\varphi\circ\phi$, where $\phi:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{n\times c}$ maps the input scan $P\in\mathbb{R}^{n\times 3}$ to hyper-dimensional features $F\in\mathbb{R}^{n\times c}$; and $\varphi:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{d}$ aggregates the local features $F$ into a global descriptor $D\in\mathbb{R}^{d}$ with $d$ dimensions.
@@ III-C1 Training Data (B1)
Place recognition methods traditionally resort to contrastive learning techniques, which learn to distinguish between similar (positive pairs) and dissimilar (negative pairs) data samples. In place recognition for urban environments, positive pairs are defined only by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in these environments. However, to perform place recognition in crops with narrow rows, the radius is a necessary but not sufficient condition. The search conditions must guarantee that the search occurs along the row and does not transgress into neighboring rows. Dividing the crops into distinct segments and confining the search within each segment ensures the necessary conditions.
@@ III-C3 Contrastive Loss (B3)
where $d_{AP}=\|D^{A}-D^{P}\|_{2}$ is the Euclidean distance between the anchor and the positive, $d_{AN}=\min\limits_{j}(\|D^{A}-\{D^{N}_{j}\}\|_{2})$ is the Euclidean distance between the anchor and the hardest negative i.e., the negative that is the closest w.r.t. $D^{A}$ in the descriptor space, while $m$ designates a margin value.
@@ IV EXPERIMENTAL EVALUATION
This section outlines the evaluation protocol, implementation details, and training procedures of the proposed approach. Both quantitative and qualitative retrieval results are presented to support the proposed contributions. Additionally, a runtime analysis is performed to assess the computational efficiency of PointNetPGAP.
@@ IV-A The HORTO-3DLM Dataset
TABLE I: Summary of the four sequences. The various sequence names represent the first letter of plantation type, month and year of the recording. For example, an orchard sequence recorded in November 2022 is named ON22. The columns have the following meaning: name of the sequence (Seq), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
@@
IV-B HORTO-3DLM Dataset for Place Recognition
@@ IV-C2 Cross-Domain
IV-D Training and Implementation Details
@@
IV-E Retrieval Evaluation and Discussion
@@
IV-E1 Ablation Study
The ablation study is conducted following the cross-validation protocol outlined in Section IV-C1, with results presented in Table V. For this analysis, we report Recall@1 and Recall@1% for each aggregation approach, both individually and combined. The results suggest that combining both aggregators results in superior performance compared to using each aggregator separately, with an improvement of 0.6 percentage points (pp) in retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate. These findings indicate that the proposed aggregator significantly enhances the model’s ability to identify the nearest neighbor.
@@ IV-E1 Ablation Study
IV-E2 Cross-Validation (same domain)
The results of this experiment are presented in Table IV. PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is further illustrated in Fig. 6, where a qualitative assessment shows the true positive predictions of each model for the Top-1 candidate retrieval along the paths. Given that PNPGAP and PointNetVLAD share the same backbone, these results suggest that the proposed aggregator is more effective in horticultural environments, particularly in handling sparse scans, than other first- and second-order aggregators.
@@ IV-E2 Cross-Validation (same domain)
However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows (i.e., 0, 1, 2, and 5). This improvement suggests that the SLC model produces more robust descriptors that are less susceptible to ambiguities.
@@ IV-E3 Cross Domain
IV-F Runtime Analysis
Processing capacity and model size are critical factors for real-world implementation, particularly in resource-constrained robotics applications. The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-D. The reported times, measured in milliseconds (ms), represent the duration required by each model to process a batch of 20 scans. Retrieval duration was excluded from this experiment, as it remains consistent across all models. Model size is expressed by the number of parameters, in millions (M). PNPGAP is notably compact, with 0.4 million parameters, and processes a batch in an average of 4.4 ms.
@@ ACKNOWLEDGMENTS
This work has been supported by the project GreenBotics (ref. PTDC/EEI-ROB/2459/2021), funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, and by Innovate UK under the Grant No. 10072930 and 10073332. It was also partially supported by FCT through grant UIDB/00048/2020 and under the PhD grant with reference 2021.06492.BD. The authors would also like to thank Dr Charles Whitfield at NIAB East Malling for facilitating orchard data collection campaigns.

### v4

@@ Abstract
Index Terms:
Deep Learning Methods; Localization; Agricultural Automation
@@
00footnotetext: This work has been supported by the project GreenBotics (ref. PTDC/EEI-ROB/2459/2021), funded by Fundação para a Ciência e a Tecnologia (FCT), Portugal, and by Innovate UK under the Grant No. 10072930 and 10073332. It was also partially supported by FCT through grant UIDB/00048/2020 and under the PhD grant with reference 2021.06492.BD.00footnotetext: 1T.Barros, L.Garrote, P.Conde, C.Premebida and U.J.Nunes are with the University of Coimbra, Institute of Systems and Robotics, Department of Electrical and Computer Engineering, Portugal. E-mails: {tiagobarros, garrote, pedro.conde,  cpremebida, urbano}@isr.uc.pt00footnotetext: 2M.J.Coombes and C.Liu are with the Dept. of Aeronautical and Automotive Engineering, LUCAS Lab, Loughborough University, UK. E-mails:{M.J.Coombes, C.Liu5}@lboro.ac.uk
@@ I INTRODUCTION
In recent years, significant advancements have been made in 3D LiDAR-based place recognition, driven primarily by the autonomous driving community with the aim of achieving long-term localization [1, 2]. Although initial efforts have been made to adapt these methods for agricultural environments [3, 4], several challenges unique to these settings remain largely unresolved. Unlike urban environments, where LiDAR scans capture well-defined geometries (e.g., corners, planes), horticultural environments produce scans with sparse and poorly defined features. The semi-permeable nature of these environments allows laser beams to traverse multiple row layers, resulting in scans that cover several rows. This phenomenon leads to sparse scans that overlap with scans from adjacent rows, thereby causing both intra- and inter-row descriptor ambiguity.
@@
We benchmark PointNetPGAP at both the sequence and segment levels against state-of-the-art (SOTA) models, specifically OverlapTransformer [5], PointNetVLAD [6], and LOGG3D-Net [7], using the HORTO-3DLM dataset within a cross-validation protocol. Additionally, we evaluate the cross-domain generalization capabilities using the KITTI Odometry dataset [8]. The results highlight that the aggregation approach of PointNetPGAP is better suited to horticultural environments compared to existing SOTA models. Although its impact is reduced in urban settings, it nonetheless exhibits superior rotation invariance in urban environments compared to SOTA models. Moreover, when models are trained with the SLC approach, results at both the sequence and segment levels indicate improved retrieval performance, particularly in segments with higher ambiguity.
@@ II RELATED WORK
In the context of 3D LiDAR-based methods, although hand-crafted descriptors are still being used [9], Deep Learning (DL) techniques have emerged as the most prevalent choice for place modeling [1]. DL-based place recognition models can be broadly split into two main components: a local feature extraction module that maps the input scan to a local feature space, and a feature aggregation module that aggregates the local features into a global descriptor.
As for the feature extraction module, approaches can be categorized based on how the point clouds are processed: those that extract features directly from point clouds, such as PointNetVLAD [6], and those that resort to proxy representations, such as voxels [10, 11], polar coordinates [12], or depth range images [13, 5]. Recent works such as MinkLoc3D [11], LCDNet [14], and LOGG3D-Net [7] tend to primarily use voxel-based approaches, employing sparse convolutions such as PV-RCNN [15] or 4D Spatio-Temporal ConvNets [16], which have shown higher feature extraction capabilities in urban-like environments. In this work, we resort to PointNet [17] due to its efficiency and lower computational demands (as demonstrated in Section IV-E).
Regarding aggregation strategies, most 3D LiDAR place recognition frameworks rely on first-order aggregation approaches such as NetVLAD [6, 18, 14], which computes the residuals w.r.t. feature clusters, generalized-mean pooling (GeM) [11] or global max pooling (MAC) [19] which operate directly on the feature space. Inspired by these works, we propose an aggregation approach that combines, in a single descriptor, a first-order and a second-order aggregation approach. Specifically, we combine the global average pooling aggregator with a pairwise feature interaction aggregator into a single descriptor. We compare our hypothesis with the SOTA models LOGG3D-Net [7], PointNetVLAD [6], and OverlapTransformer [5], and show through empirical results on real-world datasets that the combination of these two statistical-based aggregators yields higher performance than each aggregator individually and also the SOTA models.
@@ III Proposed Approach
This section introduces the PointNetPGAP and the proposed training regime, where a traditional contrastive learning approach is combined with the proposed SLC model. Figure 2 outlines both PointNetPGAP in a retrieval framework (Top) and PointNetPGAP in the proposed training scheme (Bottom).
@@ III-B PointNetPGAP
PointNetPGAP, defined here as a mapping function $\Theta$ (outlined in Fig. 2 by T1, T2), can be decomposed into two functions: $\Theta\equiv\varphi\circ\phi$, where $\phi:\mathbb{R}^{n\times 3}\rightarrow\mathbb{R}^{n\times c}$ maps the input scan $P\in\mathbb{R}^{n\times 3}$ to hyper-dimensional features $F\in\mathbb{R}^{n\times c}$ with $c$ dimensions; and $\varphi:\mathbb{R}^{n\times c}\rightarrow\mathbb{R}^{d}$ aggregates the local features $F$ into a global descriptor $D\in\mathbb{R}^{d}$ with $d$ dimensions.
@@ III-C1 Training Data (B1)
Place recognition methods traditionally rely on contrastive learning techniques, which are designed to differentiate between similar (positive pairs) and dissimilar (negative pairs) data samples. In urban environments, positive pairs are typically defined by a fixed radius, which is a sufficient condition given the scale of the trajectories and revisits in such environments. However, in narrow-row crop environments, a fixed radius is a necessary but not sufficient condition for place recognition. Additional constraints are required to ensure that loops occur within the same row, rather than between neighboring rows. To address this, crops are divided into distinct segments, and loops must not only satisfy the radius condition but also originate from the same segment.
@@ III-C3 Contrastive Loss (B3)
where $d_{AP}=\|D^{A}-D^{P}\|_{2}$ is the Euclidean distance between the anchor and the positive, $d_{AN}=\min\limits_{j}(\|D^{A}-\{D^{N}_{j}\}\|_{2})$ is the Euclidean distance between the anchor and the hardest negative: i.e., the negative that is the closest w.r.t. $D^{A}$ in the descriptor space, while $m$ designates a margin value.
@@ IV EXPERIMENTAL EVALUATION
This section outlines the evaluation protocols, implementation details, and training procedures of the proposed approach. Both quantitative and qualitative retrieval results are presented to support the proposed contributions. Additionally, a runtime analysis is performed to assess the computational efficiency of PointNetPGAP.
@@ IV-A The HORTO-3DLM Dataset
TABLE I: Summary of the four sequences. The columns have the following meaning: name of the sequence (Seq), month of recording (M), year of recording (Y), number of rows each sequence comprises (N.Rs), number of segments each sequence comprises (N.Segs), traveled distance (D) and plantation type.
@@
IV-A1 HORTO-3DLM Dataset for Place Recognition
@@ IV-B2 Cross-Domain
IV-C Training and Implementation Details
@@
IV-D Retrieval Evaluation and Discussion
@@
IV-D1 Ablation Study
The ablation study is conducted following the cross-validation protocol outlined in Section IV-B1, with results presented in Table V. For this analysis, we report Recall@1 and Recall@1% for each aggregation approach, both individually and combined. The results suggest that combining both aggregators results in superior performance compared to using each aggregator separately, with an improvement of 0.6 percentage points (pp) in retrieving the Top-1% candidates and 4.7 pp for retrieving the Top-1 candidate. These findings indicate that the proposed aggregator significantly enhances the model’s ability to identify the nearest neighbor.
@@ IV-D1 Ablation Study
IV-D2 Cross-Validation (same domain)
The results of this experiment are presented in Table IV. PNPGAP emerges as the best model when compared to the SOTA models, outperforming PointNetVLAD by 13.6 pp without the SLC model and by 12.4 pp with the SLC model when retrieving the Top-1 candidate.This superiority is illustrated further in Fig. 6, where a qualitative assessment shows the true positive predictions of each model for the Top-1 candidate retrieval along the paths. Given that PNPGAP and PointNetVLAD share the same backbone, these results suggest that the proposed aggregator is more effective in horticultural environments, particularly in handling sparse scans, than other first- and second-order aggregators.
@@ IV-D2 Cross-Validation (same domain)
However, when the SLC model is incorporated, the results at the segment level indicate a notable performance increase, particularly in the segments corresponding to rows. This improvement suggests that the SLC model produces more robust descriptors that are less susceptible to ambiguities.
@@ IV-D3 Cross-Domain
IV-E Runtime Analysis
Processing capacity and model size are critical factors for real-world implementation, particularly in resource-constrained robotics applications. The runtime experiment results, presented in Table VII, are obtained using the hardware detailed in Section IV-C. The reported times, measured in milliseconds (ms), represent the duration each model requires to process a batch of 20 scans. Retrieval duration was excluded from this experiment, as it remains consistent across all models. Model size is expressed by the number of parameters, in millions (M). PNPGAP is notably compact, with 0.4 million parameters, and processes a batch in an average of 4.4 ms.
@@ ACKNOWLEDGMENTS
The authors would also like to thank Dr Charles Whitfield at NIAB East Malling for facilitating orchard data collection campaigns.